OpenStudy (anonymous):

The table lists data regarding the average salaries of several professional athletes in the years 1991 and 2001. a) Use the data points to find a linear function that fits the data . b) Use the function to predict the average salary in 2005 and 2010. Year 1991 the average salary is 269,000 year 2010 the average salary is 1,390,000

OpenStudy (anonymous):

Treat this as a linear equation (y=mx+b). The year is your x value, and the salary is your y-value, so you really have two ordered pairs - (1991, 269000) and (2010, 1390000). Use these to find the slope (m) of your line.

OpenStudy (anonymous):

Once you have the slope, then use the values from one of the ordered pairs to find the y-intercept (b).

OpenStudy (anonymous):

I need to solve the linear function that fits data S(x)= And then predicted the average salary for 2005 and 2010?

OpenStudy (anonymous):

Is the data you listed above all of the data that is given?

OpenStudy (anonymous):

yes sir!

OpenStudy (anonymous):

Then what I said above should work, though your equation will take the form S(x)=mx+b, since it is supposed to be a function. Use the ordered pairs to find the slope (change in y value over change in x value).

OpenStudy (anonymous):

Plug the two points into the slope formula to find the slope of the linear function. Then plug the slope and one of the points into the point slope formula: \[y-y_1 = m(x-x_1)\] Where \((x_1,y_1)\) is one of your points and m is the slope.

OpenStudy (anonymous):

That make no scence to me at all.....!

OpenStudy (anonymous):

Ok, do you know the formula for the slope if you have 2 points?

OpenStudy (anonymous):

x1 and y2 over x2 and y2 but it is hard for to get the answer

OpenStudy (anonymous):

That's not quite right. \[\frac{y_2-y_1}{x_2-x1} = Slope\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So plug in the salaries as y2 and y1 and the years as x2 and x1. Make sure you keep them consistent, so the salary for one year is y2 then that year must be x2.

OpenStudy (anonymous):

And then post what you get for slope

OpenStudy (anonymous):

11,210,000 over 19

OpenStudy (anonymous):

You have an extra 0 there. Should be 1,121,000 over 19

OpenStudy (anonymous):

Ok, so now plug in that slope, along with one of your two points into the point slope formula.

OpenStudy (anonymous):

yea sry so the answer is 1,121,000 over 19 is s(x)?

OpenStudy (anonymous):

No, that's the slope of your line.

OpenStudy (anonymous):

Now you plug in that slope along with one of your points into the point slope formula: \[y-y_1 = m(x-x_1)\] Where \(x_1,y_1\) are the x and y values of your point, and m is the slope.

OpenStudy (anonymous):

I am lost!!!! SRY

OpenStudy (anonymous):

You have the slope. The slope is m. you have 2 points. Pick one.

OpenStudy (anonymous):

I am not good at this type of question

OpenStudy (anonymous):

Which point do you want to use?

OpenStudy (anonymous):

269,000

OpenStudy (anonymous):

That's a salary, the point would be (1991,269,000)

OpenStudy (anonymous):

1991

OpenStudy (anonymous):

Where 1991 is your \(x_1\) and 269,000 is your \(y_1\). Plug them into the equation and plug in the slope you found for m. That will give you the equation for the line which represents the salary over time.

OpenStudy (anonymous):

so what is s(x)= then

OpenStudy (anonymous):

After you plug it in, s(x) = y

OpenStudy (anonymous):

I also have to predict the salary for 2005 and 2010?

OpenStudy (anonymous):

Which you can do easily once you have the formula. Have you plugged it in yet?

OpenStudy (anonymous):

so you have 1,120,000/9=124,444

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

1,120,000/19=58,947

OpenStudy (anonymous):

I am so confused!!!!

OpenStudy (anonymous):

You picked your point, so you should have \[y - 269,000 = \frac{1,121,000}{19}(x - 1991)\]

OpenStudy (anonymous):

well 1,121,000 / 19=59,000

OpenStudy (anonymous):

Solving for s(x) = y we have: \[ s(x) = y = \frac{1}{19}(1,121,000x -2226800000)\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So that's the equation for the salary at a given year x.

OpenStudy (anonymous):

You want to know the salary at 2010, plug that in for x, and see what y is.

OpenStudy (anonymous):

or s(x), whichever you prefer.

OpenStudy (anonymous):

2,030,530,000

OpenStudy (anonymous):

26,410,000 to many zero when working

OpenStudy (anonymous):

2010 would be 26,410,000 and 2005 would be 20,805,000

OpenStudy (anonymous):

I am sorry if I am causing you grief I am just bad at these problem.